### v1

# A quantitative analysis of Galilei’s observations of Jupiter satellites from the Sidereus Nuncius

Abstract
We analyse the observations of the satellites of Jupiter from the Sidereus Nuncius (January 7 to March 1, 1610) and compare them to the predictions obtained using a modern sky simulator, verifying them one by one. A sinusoidal fit of the data obtained from the 64 available sketches, allows measuring the relative major semi-axes of the satellites’ orbits and their periods with a statistical precision of 2-4% and 0.1-0.3% respectively. The periods are basically unbiased while the orbits tend to be underestimated for Callisto by about 12%. The posterior fit error indicates that the positions of the satellites are determined with a resolution of 0.4-0.6 Jupiter diameters in the notation of Galilei corresponding to about 40-70 arc sec i.e. similar to the true angular diameter of Jupiter, in those days. We show that with this data one can infer in a convincing way the third law of Kepler for the Jupiter system. The 1:2 and 1:4 orbital resonance between the periods of Io and Europa/Ganymede can be determined with % precision. In order to obtain these results it is important to separate the four datasets. This operation, which is nowadays simple using a sky simulator, and is fully reported in this work, was an extremely difficult task for Galilei as the analysis will evidence. Nevertheless we show how the four periods might have been extracted using the modern Lomb-Scargle technique without having to separate the four data-sets already just using these early observations. We also perform a critical evaluation of the accuracy of the observation of the Pleiades and other clusters and the Moon. To conclude we use these data to infer the possible resolution and aberration of the “cannocchiale” of Galilei and use these parameters to imagine what could have been the view offered by Galilei to the Doge when he was promoting his instrument from the Campanile of San Marco in Venice. This analysis allows quantifying and appreciating the level of accuracy that Galilei managed to attain with the instrumentation he had at his disposal.
keywords:
Galilei, Sidereus Nuncius, Galilean moons, reanalysis
1 Introduction
The night of January 7 1610 was a special one for Galilei and, in hindsight, for the history of science. With an improved version [^1] Interestingly, Galilei writes that he had not seen the new stars with the previous version of the telescope. of his telescope, that he had already advertised about six months earlier to the headquarters of the Venetian republic, Galilei points to Jupiter.
He sees three peculiar brilliant “stars” very close to the planet and strangely well aligned to the Ecliptic plane (not a random one!). See Fig. 22. Probably by false modesty, he writes he had not paid too much attention to the relative distances as he expected them to be fixed stars and, guided by I know not what fate, (nescio quo fato ducto) he came back to observe it the night after. It is hard to believe he did not already had an intuition of the exceptional discovery that was behind the corner. Which is the probability of having by change three aligned stars so closely packed to Jupiter? And in that very meaningful direction? Quite small indeed. But something even more amazing was happening the day after.
On January 8 Jupiter appears to the East of the three stars (Fig. 23). In reality the easternmost “star” (Callisto) was still there to the East of Jupiter, but Galilei does not see it, unlikely to the earlier night. It is interesting to notice that Callisto will always be observed in the following two months, unless when it was too close to Jupiter. The three stars were then in reality not the same three satellites: Io had passed from the East to the West side and Europa had become visible to the West after being too close to Jupiter the day before.
If the stars were fixed the explanation could be that Jupiter was moving in the “wrong” direction (eastward) in the sky. Galilei thinks that Jupiter, as well as Mars, might be travelling in retrograde motion. Clouds on the 9th create some frustration but a new observation becomes possible on the 10th: Jupiter is now West of two stars now!
A thrilling adventure that will accompany him uninterruptedly for each clear night for the following two months was just about to start [^2] It is remarkable how many clear days there were in Padua during those winter months: just ten days were cloudy out of about sixty. On average nowadays the fraction of cloudy days is 40-50% according to [1].. This detailed account of the dance of the satellites of Jupiter is an history of scientific rigour and excitement: it lead him to observe until late at night in several occasions despite the many duties that he must have had in Padua.
Earlier in the book Galilei writes that astronomers should engage in determining the periods of the satellites as this was not possible to him due to the restrictions in time. In view of this consideration it is amazing how accurately Galilei decided to document this early phenomenon. Despite his statement these 64 observations were indeed extremely rich and go far beyond what one would deem sufficient to state a new “discovery”.
It is possible that these observations seemed so exceptional and groundbreaking to him to deserve such a solid and complete basis of evidence despite this might seem redundant to our eyes.
It should also be not forgotten that Galilei might have felt the pressure of releasing the book as soon as possible due to the risk of being anticipated by other scientists, which was not completely an unlikely scenario. Despite this he deemed necessary to report 64 observations. This choice tells a lot about of the mindset of Galilei and this emphasis on data cannot but induce a certain degree of admiration in modern experimentalists.
In this work we analyse the observations of the satellites of Jupiter reported by Galileo Galilei in the Sidereus Nuncius [2] (“The stellar messenger”) during the period from January 7 to March 1, 1610. The 64 sketches that Galilei presented in his revolutionary work are presented in Sec. 2 and the result of the image digitization in Sec. 3. In Sec. 3.1 we convert the time of the observations from “Italian hours”, calculated starting from sunset, into modern hours (CET) to match it with the predictions of an online sky simulator [3] “Stellarium” (Sec. 3.4). Before performing the association of the measurements to the individual four satellites, we discuss the possibility of determining the orbital parameters from a global analysis of all the points taken together (Sec. 3.3): we consider the statistical distribution of the satellites position and an analysis based on the Lomb-Scargle periodogram in Sec. 3.3.
While performing the association of the satellites (Sec. 3.4)) we highlight in detail the possible ambiguities due to the fact that sometimes a pair of close satellites were reported as that they could be missed due to their proximity to the planet’s disk. We also show how, generally, the observation have a clear mapping to the expectation from the simulation.
A sinusoidal fit of the data obtained from the data-sets divided by satellite, is presented in Sec. 3.6. We discuss the achievable accuracy on the determination of the major semi-axes of the satellites’ orbits and their periods and the residuals between the observations and the fit. We also consider the origin of the observed biases with respect to modern determinations. We show that with this data one can infer in a convincing way the equivalent of the third law of Kepler for the Jupiter system. We show that the 1:2 and 1:4 orbital resonances between the periods of Io : Europa and Io : Ganymede can be determined with % precision from these data.
In Sec. 4 we expand the study by performing a critical evaluation of the accuracy of the observation of the Pleiades cluster, the Beehive cluster, the region of Orion belts, what Galilei called the ”head of Orion Nebula” and finally the Moon.
To conclude in Sec. 3.7 we use the accuracy of Jupiter data to infer the possible resolution and aberration of the “cannocchiale” of Galilei. We use these parameters to imagine what could have been the view offered by Galilei to the head of the republic of Venice (the “Doge” of the “Serenissima”) when he was promoting his instrument from the Campanile of San Marco in Venice.
2 Available dataset from the “Stellar messenger”
In this work we decided to only focus on the observations presented in the Sidereus Nuncius without considering for simplicity later observations performed in later years by Galileo himself. The images were taken from [2]. We show in Figs. 1-2 a collection of the sketches of Sidereus Nuncius in chronological order with a progressive numeric labeling.
1)Refer to caption
2)Refer to caption
3)Refer to caption
4)Refer to caption
5)Refer to caption
6)Refer to caption
7)Refer to caption
8)Refer to caption
9)Refer to caption
10)Refer to caption
11)Refer to caption
12)Refer to caption
13)Refer to caption
14)Refer to caption
15)Refer to caption
16)Refer to caption
17)Refer to caption
18)Refer to caption
19)Refer to caption
20)Refer to caption
21)Refer to caption
22)Refer to caption
23)Refer to caption
24)Refer to caption
25)Refer to caption
26)Refer to caption
27)Refer to caption
28)Refer to caption
29)Refer to caption
30)Refer to caption
31)Refer to caption
32)Refer to caption
Figure 1: Sidereus Nuncius sketches A, from January 7 to February 1st.
33)Refer to caption
34)Refer to caption
35)Refer to caption
36)Refer to caption
37)Refer to caption
38)Refer to caption
39)Refer to caption
40)Refer to caption
41)Refer to caption
42)Refer to caption
43)Refer to caption
44)Refer to caption
45)Refer to caption
46)Refer to caption
47)Refer to caption
48)Refer to caption
49)Refer to caption
50)Refer to caption
51)Refer to caption
52)Refer to caption
53)Refer to caption
54)Refer to caption
55)Refer to caption
56)Refer to caption
57)Refer to caption
58)Refer to caption
59)Refer to caption
60)Refer to caption
61)Refer to caption
62)Refer to caption
63)Refer to caption
64)Refer to caption
Figure 2: Sidereus Nuncius sketches B, from February 2 to March 1st.
In total Galileo took 64 sketches in 1610 from January 7th up to March 1st basically inspecting Jupiter whenever the meteorological conditions in Padua allowed him. As it became evident that the change in the configurations had visible changes on a scale of a few hours [^3] The period of Io is just about 1.8 days. Galileo decided to observed the system at different times during the same night. Such multiple observations occurred for example on Jan. 15 (id 7, 8), Jan. 17 (id 10, 11), Jan. 19 (id 13, 14), Jan. 20 (id 15, 16, 17), Jan. 22 (id 19, 20), Jan. 23 (id 21, 22), Jan. 24 (id 23, 24), Jan. 26 (id 26, 27), Jan. 31 (id 30, 31), Feb. 2 (id 33, 34), Feb. 4 (id 36, 37), Feb. 11 (id 43, 44, 45), Feb. 15 (id 48, 49, 50), Feb. 18 (id 53, 54), Feb. 26 (id 58, 59), Feb. 28 (id 62, 63). The range of these measurements taken during the same night spans up to about six hours. On Jan. 20 for example the first observation was done 1h15’ after sunset and the last 7h after sunset so at about 6 p.m. and 1 a.m.. Generally Galilei reported the configuration of the satellites only except in four occasions (observation 60, 61, 62, 64) when he also noted down the position of a nearby star [^4] We have identified this star as HD32811 or SAO 76962m or HIP 23784. It has mag. 7.16, a distance of 872.09 light years, type B8/9V and is located at 04h 43m 18.2s +21o549.9” according to the software Stellarium.. This was intended to clarify the motion of Jupiter and of its satellites with respect to the “really-fixed” stars. The satellites are marked as asterisks and the size of the symbol is a proxy for the relative brightness. In many occasions Galileo also documented the fact that the satellites appeared to be not completely aligned (see for example observations 6, 7, 11, 16, 17, 19, 23, 27, 30, 31). As we will see in the following when comparing with modern simulation also the recordings of these displacements from the ecliptic plane are remarkably accurate.
3 Digitization of the sketches
Tables 1 and 2 summarize the measurements; the first table refers to the observations performed in January while the second to those of February and March. Data have been obtained from the analysis of digitized images using cursors that were manually centred on the satellites. For each image the size of the planet disk was also recorded to normalize the elongations in terms of this diameter. The diameter has always been found to be constant between the sketches with few exceptions (60, 61, 62, 64) where also the nearby star is drawn. The first column is a sequential identifier that correspond to the one of Fig. 1. The following three columns give the time and hour of the observation as reported by Galilei (in italic hours). The 5th column reports the number of visible satellites. The last columns show the position of the satellites ($x_{i}$, $y_{i}$) where $x$ runs horizontally from left to right and $y$ vertically from bottom to top. In the sketches the East is left and West right. This is the most natural choice but is also motivated by the fact that Galilei’s telescope was not inverting the image being composed of a semi-convex or convex objective and bi-concave or plano-concave eyepiece.
id m day h:min $n$ $(x_{0},y_{0})$ $(x_{1},y_{1})$ $(x_{2},y_{2})$ $(x_{3},y_{3})$
1 1 7 1:00 3 (-5.72, 0) (-2.28, 0) (2.42, 0)
2 1 8 / 3 (1.53, 0) (3.14, 0) (4.78, 0)
3 1 10 / 2 (-4.09, 0) (-1.98, 0)
4 1 11 / 2 (-4.95, 0) (-3.64, 0)
5 1 12 1:00 3 (-2.73, 0) (-1.36, 0) (1.50, 0)
6 1 13 / 4 (-2.16, 0) (0.95, 0) (1.64, 0.11) (2.25, 0)
7 1 15 3:00 4 (2.05, 0) (3.41, 0) (4.77, 0) (7.05, -0.52)
8 1 15 7:00 3 (3.14, 0) (4.09, 0) (6.43, 0)
9 1 16 1:00 3 (-0.98, 0) (1.02, 0) (5.41, 0)
10 1 17 0:30 2 (-1.95, 0) (4.77, 0)
11 1 17 5:00 3 (-2.82, 0.05) (-2.25, -0.05) (5.95, -0.14)
12 1 18 0:20 2 (-4.50, 0.09) (5.95, -0.18)
13 1 19 2:00 3 (-3.50, 0) (3.11, 0) (5.43, 0)
14 1 19 5:00 4 (-3.32, 0) (-1.48, 0) (2.75, 0) (4.39, 0)
15 1 20 1:15 3 (-1.00, 0) (0.95, 0) (1.75, 0)
16 1 20 6:00 3 (-1.59, 0) (1.14, 0.18) (1.73, 0.18)
17 1 20 7:00 4 (-2.00, 0) (0.82, 0.14) (1.41, 0) (1.82, 0.14)
18 1 21 0:30 4 (-1.95, 0) (-1.45, 0) (-1.09, 0) (2.39, 0)
19 1 22 2:00 4 (-3.36, 0) (1.32, 0) (1.93, -0.09) (4.11, -0.09)
20 1 22 6:00 4 (-3.32, 0) (1.23, 0) (2.09, 0) (2.84, 0)
21 1 23 0:40 3 (-4.86, 0) (-1.82, 0) (1.89, 0)
22 1 23 5:00 1 (-4.27, 0)
23 1 24 / 3 (-6.82, 0) (-2.95, -0.14) (-2.14, 0)
24 1 24 6:00 2 (-5.27, 0) (-1.59, 0)
25 1 25 1:40 2 (-6.82, 0) (-3.68, 0)
26 1 26 0:40 3 (-6.77, 0) (-3.36, 0) (3.43, 0)
27 1 26 5:00 4 (-7.32, 0) (-3.70, 0) (-1.75, 0.18) (3.45, 0)
28 1 27 1:00 1 (-4.91, 0)
29 1 30 1:00 3 (-1.95, 0) (2.30, 0) (3.07, 0)
30 1 31 2:00 3 (-2.93, 0) (-2.23, 0.18) (4.80, 0)
31 1 31 4:00 3 (-2.32, 0) (-1.59, 0.20) (5.75, 0)
Table 1: Dataset 1 (January). The last four columns contain the position of the satellites from left to right in units of the Jupiter equatorial diameter, as in the sketches of Galilei.
id m day h:min $n$ $(x_{0},y_{0})$ $(x_{1},y_{1})$ $(x_{2},y_{2})$ $(x_{3},y_{3})$
32 2 1 2:00 3 (-4.66, 0) (-1.00, 0) (4.77, 0)
33 2 2 / 3 (-4.30, 0) (2.48, 0) (6.73, 0.14)
34 2 2 7:00 4 (-4.39, -0.05) (-1.95, 0) (3.36, 0.05) (7.27, 0.14)
35 2 3 7:00 3 (-1.55, 0) (1.68, 0) (7.00, 0.18)
36 2 4 2:00 4 (-2.86, 0) (-1.18, 0) (2.64, 0) (5.64, 0)
37 2 4 7:00 4 (-2.50, 0) (-1.77, 0) (2.59, 0) (4.32, 0)
38 2 6 / 2 (-2.14, 0) (2.34, 0)
39 2 7 / 2 (-1.95, 0) (-1.18, 0)
40 2 8 1:00 3 (-4.36, 0) (-3.64, 0) (-1.55, 0)
41 2 9 0:30 3 (-6.45, 0) (-4.25, 0) (2.86, 0)
42 2 10 1:30 2 (-6.18, 0.14) (-1.02, 0)
43 2 11 1:00 3 (-7.25, 0.14) (-2.50, 0) (2.84, 0)
44 2 11 3:00 4 (-6.89, 0.14) (-2.50, 0.05) (-0.95, 0) (3.07, 0)
45 2 11 5:30 4 (-7.23, 0.14) (-2.59, 0) (-1.55, 0) (2.84, 0)
46 2 12 0:40 4 (-6.09, 0.14) (-0.82, 0) (1.09, 0) (5.32, -0.14)
47 2 13 0:30 4 (-4.82, 0) (-1.95, 0) (2.45, 0) (3.20, 0)
48 2 15 / 3 (-2.86, 0) (-1.55, 0) (-1.14, 0)
49 2 15 5:00 2 (-2.89, 0) (-0.91, 0)
50 2 15 6:00 3 (-3.07, 0) (-0.98, 0) (1.86, 0)
51 2 16 6:00 3 (-4.77, -0.16) (3.45, 0.14) (5.09, 0.18)
52 2 17 1:00 2 (-2.68, 0) (6.18, 0.14)
53 2 18 / 3 (-2.50, 0) (1.70, 0.09) (6.11, 0.23)
54 2 18 6:00 4 (-3.07, 0) (-1.36, 0) (2.48, 0) (7.27, 0.23)
55 2 19 0:40 2 (4.00, 0) (7.23, 0)
56 2 21 1:30 3 (-2.14, 0) (2.07, 0) (6.09, 0)
57 2 25 1:30 3 (-4.77, 0) (-2.68, 0) (1.89, 0)
58 2 26 0:30 2 (-5.73, 0) (4.16, 0)
59 2 26 5:00 3 (-6.09, 0) (1.36, 0) (4.20, 0)
60 2 26 5:00 4 (-6.53, 0) (0.94, 0) (4.18, 0)
61 2 27 1:04 4 (-5.05, 0) (-0.70, 0) (1.68, 0) (2.43, 0)
62 2 28 1:00 2 (1.74, 0)
63 2 28 5:00 3 (-6.364, 0) (-1.545, 0) (1.682, 0)
64 3 1 / 4 (-5.18, 0) (-3.41, 0.24) (-2.53, 0) (-1.71, 0)
Table 2: Dataset 2 (February-March). The last four columns contain the position of the satellites from left to right in units of the Jupiter equatorial diameter, as in the sketches of Galilei.
A graphical summary of the digitized positions of the satellites, normalized to the diameter of Jupiter as in the previous tables is shown in Fig. 3. The time flows from top to bottom. Comparing this plot with Figs. 1 and 2 ensures the absence of mistakes in the digitization process.
Figure 3: Reconstructed positions of Jupiter satellites from digitization of the sketches in units of the planet’s equatorial diameter. Times increases from top to bottom.
3.1 Conversion of the observation times
Galilei expressed the observation times in terms of hours since sunset following italic hours that was used at those times. The first hour, for example, refers to one hour after sunset. To convert these data into modern hours (CET) we considered the time of sunset in Padua in the beginning of the year (see Fig. 86, Appendix) and sum the hour reported by Galilei to this value assuming that each hour has the same length. The time of sunset was approximate linearly as: $h_{sunset}=16.58+0.02353~{}d$ where hours are expressed in decimal fractions and $d$ is the ordinal day since the beginning of the year. The result of the conversion is shown in Tab. 3 and 4 where all the observations are associated to a time expressed in CET. It should be noted that 11 times Galilei does not mention the time of the observations (marked as “ïn the column 4 Tab.3 and 4). In this case for the subsequent analysis we assumed that the time was the first hour, a frequent choice. This introduces for these observations an error up to potentially 5-6 hours [^5] The latest observations are before 1 AM. Note that the setting time of Jupiter was about 5.30 and 2.00 at the beginning and end of the considered period. which can be relevant for satellites passing close to Jupiter.
id month day after sunset (h) sunset modern time (h:min)
1 1 7 1:00 16.74 17:44
2 1 8 1:00 16.77 17:46
3 1 10 1:00 16.81 17:48
4 1 11 1:00 16.84 17:50
5 1 12 1:00 16.86 17:51
6 1 13 1:00 16.89 17:53
7 1 15 3:00 16.93 19:55
8 1 15 7:00 16.93 23:55
9 1 16 1:00 16.96 17:57
10 1 17 0:30 16.98 17:28
11 1 17 5:00 16.98 21:58
12 1 18 0:20 17.00 17:20
13 1 19 2:00 17.03 19:01
14 1 19 5:00 17.03 22:01
15 1 20 1:15 17.05 18:17
16 1 20 6:00 17.05 23:02
17 1 20 7:00 17.05 24:02
18 1 21 0:30 17.07 17:34
19 1 22 2:00 17.10 19:05
20 1 22 6:00 17.10 23:05
21 1 23 0:40 17.12 17:47
22 1 23 5:00 17.12 22:07
23 1 24 1:00 17.14 18:08
24 1 24 6:00 17.14 23:08
25 1 25 1:40 17.17 18:50
26 1 26 0:40 17.19 17:51
27 1 26 5:00 17.19 22:11
28 1 27 1:00 17.21 18:12
29 1 30 1:00 17.29 18:17
30 1 31 2:00 17.31 19:18
31 1 31 4:00 17.31 21:18
Table 3: Conversion from Italic to modern hours (January 1610).
id month day after sunset (h) sunset modern time (h:min)
32 2 1 2:00 17.33 19:19
33 2 2 1:00 17.36 18:21
34 2 2 7:00 17.36 24:21
35 2 3 7:00 17.38 24:22
36 2 4 2:00 17.40 19:24
37 2 4 7:00 17.40 24:24
38 2 6 1:00 17.45 18:26
39 2 7 1:00 17.47 18:28
40 2 8 1:00 17.50 18:29
41 2 9 0:30 17.52 18:01
42 2 10 1:30 17.54 19:02
43 2 11 1:00 17.57 18:34
44 2 11 3:00 17.57 20:34
45 2 11 5:30 17.57 23:04
46 2 12 0:40 17.59 18:15
47 2 13 0:30 17.61 18:06
48 2 15 1:00 17.66 18:39
49 2 15 5:00 17.66 22:39
50 2 15 6:00 17.66 23:39
51 2 16 6:00 17.68 23:41
52 2 17 1:00 17.71 18:42
53 2 18 1:00 17.73 18:43
54 2 18 6:00 17.73 23:43
55 2 19 0:40 17.76 18:25
56 2 21 1:30 17.80 19:18
57 2 25 1:30 17.90 19:23
58 2 26 0:30 17.92 18:25
59 2 26 5:00 17.92 22:55
60 2 26 5:00 17.92 22:55
61 2 27 1:04 17.94 19:00
62 2 28 1:00 17.97 18:58
63 2 28 5:00 17.97 22:58
64 3 1 0:40 17.99 18:39
Table 4: Conversion from Italic to modern hours (February, March 1610).
3.2 Satellite-untagged analysis
Before trying to associate the observations to individual satellites, we perform a global analysis putting all the measurements together. The disentangling of the contributions of single satellites was not obvious for Galilei as the data were taken sparsely and often satellites were too close and unresolved. The oscillatory pattern of Callisto is nevertheless quite easily separable (see Fig.  3).
We first check if the distribution is consistent with the sum of four harmonic oscillators corresponding to four distinct circular orbits and then consider how the four frequencies might emerge using the Lomb-Scargle algorithm ([4, 5]).
The distribution of the horizontal displacements ($x_{i}$) is shown in Fig. 4 for the data (bullets) and the expectation in the hypothesis that the motion is harmonic. As expected the probability of observing the satellite when it is at the maximal elongation is higher as its projected velocity is smaller. Galileo did not arrive to notice that the satellites were moving faster when they were closer to Jupiter which is something that might have been at hand with his data.
Figure 4: Distribution of the elongation of the satellites for data (bullets) and a simulation assuming harmonic motions with (green) and without (magenta) the inefficiency effect in the data related to the proximity to Jupiter’s disk.
The expectation in the solid magenta line assumes that satellites are only invisible if they are behind or in front of the disk of Jupiter. The expectation is solid green line shows instead the distribution assuming that the inefficient angular region close to Jupiter corresponds to 2.5 times the real dimension of the disk. It is visible how this second hypothesis is quite consistent with the data.
The superposition of the distribution (that are each normalized to each other) was done by eye by scaling the $x$-axis to match the most extreme observations of Callisto. The dimensions of the orbits are based on modern data. It should be noted that this is very approximate as it assumes that data were taken approximately uniformly in time while, as we saw, Galilei was often taking several, near in time, measurements per night. The large statistical error allows to neglect this and we can say that the observation are semi-qualitatively consistent with what we expect. We will perform fully quantitative considerations anyway in the following.
To isolate the outermost satellite, Callisto, one could only select data being more than four diameters away from Jupiter (see Fig. 4) where only Callisto can reach. After fitting these data with a sinusoid, the points compatible with the fitted Callisto trajectory could be then removed in the innermost part. There one would have been left with three contributions and iterate. Yet looking at the data it is clear that this was not at all straightforward.
3.3 Lomb-Scargle periodograms
We now try to use the modern technique proposed by Lomb-Scargle [4, 5] to single out the emergence of the expected four distinct mono-chromatic frequencies in the data.
We show in Fig. 5 the LS periodogram of the full dataset.
Figure 5: Lomb-Scargle periodogram of all the data. Top: the time-series. Bottom: the periodogram. The coloured vertical lines denote the position where peaks are expected to occur for Callisto, Ganymede, Europa and Io, from left to right.
The top plot shows the time-series ($x_{i}$ vs $t_{i}$, with $t_{i}$ in fraction of years since 1/1/1610) while at the bottom show the corresponding LS periodogram. The python scipy.signal package was used (lombscargle). Vertical coloured lines indicate the value of the expected frequencies based on modern data. From left to right, blue is Callisto, green Ganymede, light green Europa and red, Io. It can be seen that peaking structures in the periodogram actually emerge at the correct locations. There is also a spurious one at about 1000 rad/year with a significance similar to the one of the satellites. The peak corresponding to Callisto, is, as expected, the most prominent and with the best signal-to-noise ratio while the other three have a weaker evidence. We will later re-apply the algorithm to the satellite-tagged four datasets that will be described in the next section.
3.4 Association of the satellites using a sky simulator
This association allows inspecting the expected pattern of satellites using the Stellarium sky simulator.
An example of such an exercise is shown in Fig. 6 for the observation number 7, performed on Jan. 15 1610 at the third hour (Galilei’s time) or 19.55 CET. The correspondence is remarkable and this happens for the majority of the cases as it will be shown later (see Appendix A). It is worth noticing also the level of attention of the recordings where the slight displacement of Callisto from the ecliptic plane is reported.
It is interesting to notice that, according to the simulator, the satellites’ magnitudes were, in decreasing brightness: Ganymede (4.75), Io (5.11), Europa (5.43) and Callisto (5.69). Galilei reports instead in decreasing brightness Callisto, Ganymede, Europa, Io. It is possible that the closeness to Jupiter might have played a role in biasing the perception of the brightness [^6] It should also be noted that the accuracy of the predictions of Stellarium have not been cross-checked with other simulators..
Another evident feature is the fact that the size of Jupiter’s disk is generally overestimated by Galilei. This is probably due to the limited resolution of its telescope. This interpretation is corroborated by the fact the often satellites were missed when too close to the planetary disk, as it will be detailed later.
The typographic symbol employed to denote Jupiter is circular but the inner white part seems to hint at the fact the indeed the disk has some equatorial bulging. On the other hand it should be noted that in those days Jupiter was far from opposition so the left limb was in shadow (see for example Fig. 6) making it more circular-shaped. Furthermore it should be noticed that the equatorial bulging, is a 7% effect, so it needs an appropriate resolution to be discernible.
Another feature that is apparent already from this single image is that the positions of the outer satellites Ganymede and Callisto seems to be underestimated by Galilei.
Figure 6: Observation 7, 15 January 1610, 19.55 CET: Galilei (top) vs simulation (bottom).
Each sketch was put side by side with the output of the Stellarium web-based sky simulator after taking care to see the values to those that were translated in modern time. All the images from Stellarium are shown together with the sketches of Galileo in Appendix A.
The output of the associations is given in Tabs. 5 and 6. In several cases two satellites were too close to be separated by Galileo. When this happens we decided to perform the association with one of the two. In general we find an almost perfect correspondence between the notes of Galileo and the simulator.
id comment pattern not seen
1 CI $\star$ G
2 C unseen  $\star$ IEG C
3 G-E unresolved. I too close to J CE $\star$ I
4 IE too close to J GC $\star$
5 I-C unresolved GC $\star$ E
6 $\star$ GIC
7  $\star$ IEGC
8 EGC
9 G-E unresolved $\star$ EC
10 G-E unresolved. I too close to J $\star$ C I
11 I behind GE $\star$ C
12 I-E close in transit $\star$ C
13 I unresolved too close $\star$ EC I
14 GI $\star$ EC
15 G unresolved too close at right $\star$ IC G
16 C/G unresolved $\star$ CI
17 $\star$ CGI
18 EIC $\star$ G
19 $\star$ IEG
20 $\star$ IEG
21 G/E unresolved CI $\star$ E
22 IEG too close to J $\star$ IEG
23 I too close to J on the right CEG $\star$ I
24 E-G unresolved, I behind CG $\star$
25 IE too close to J (right) CG $\star$ IE
26 I behind J CG $\star$ E
27 CGI $\star$ E
28 EI too close to J, G in transit $\star$ EI
29 G-E unresolved $\star$ GC
30 I (right) too close to J EG $\star$ C I
31 ?
32 E-I unresolved GE $\star$ C
Table 5: Summary of satellites association with comments.
id comment pattern not seen
33 $\star$ EC
34 GI $\star$ EC
35 G too close to J (right) $\star$ IC G
36 EI $\star$ GC
37 EI $\star$ GC
38 C too close to J (right), E-G unresolved $\star$ G C
39 C-E unresolved, I too close to J EG $\star$ I
40 I-E unresolved CGE $\star$
41 I too close to J (right) CG $\star$ E I
42 I-E on transit CG $\star$
43 I too close to J (left) CE $\star$ G I
44 CEI $\star$ G
45 CEI $\star$ G
46 I seen on the other side than expected (see text) CI $\star$ EG
47 CI $\star$ EG
48 C (right) too close to J GEI $\star$ C
49 I too close to J (left), C close (right) GE $\star$ IC
50 I too close (left) GE $\star$ C I
51 I behind $\star$ EC
52 I too close (left), E too close (right) $\star$ C IE
53 I too close (left) $\star$ GC I
54 EI $\star$ GC
55 I too close to J (right), E in transit  $\star$ GC I
56 G behind $\star$ IC
57 I behind CE $\star$ G
58 E (left) and I (right) too close to J $\star$ G EI
59 E too close to J (right) $\star$ IG E
60 E too close to J (right) $\star$ IG E
61 CI $\star$ EG
62 E (left) and G (right) too close to J $\star$ I EG
[…]
64 CGEI $\star$
Table 6: Summary of satellites association with comments.
In total the satellites that were potentially visible but are not recorded are 33 (Io 17 times, Europa 8 times, Ganymede 4 times and Callisto 4 times). We have excluded the cases in which the satellites were transiting the disk of Jupiter or were behind it. The reason for the inefficiency of the recordings is almost always due to the proximity of the satellite to the disk of the planet that must have been blurry. We can estimate that Galileo could see the satellites is they were away from the limb of the planet by at least approximately 2.5 diameters or about 1’40”. It should be noted that on January 7 1610 Jupiter was 4.28 AU away corresponding to about 45” for the disk angular diameter while on March 1 it was 5.02 AU away corresponding to 38”. This change in angular diameter of the system by 18% is not taken into account in the following analysis as we always report the elongations of satellites normalised to the angular diameter of the disk. Indeed the result suppor the validity of this assumption.
Observation 46 is probably one of the only that is difficult to reconcile with the prediction of the simulator. Galilei sees two satellites on the left and two on the right. It seems that either Io or Europa are on opposite sides of the disk while the simulator predicts both being on the right. We have checked if this could be related to a mismatch in the recorded hour but in the night of Feb 11 the pattern was the same as observation 45 (both Io and Europa left). The transit of both happens almost parallelly during the daytime of Feb 12 and during the early night of 12 they are both on the right side.
For observation 49 it is interesting to note that Callisto is not recorded even if it is less close than Europa that was seen instead.
3.5 Periodograms of the satellite-tagged datasets
The periodograms of the time series separated by satellite using the sky simulator are shown in Fig. 7.
Figure 7: Lomb-Scargle analysis for the Io (top left), Europa (top right), Ganymede (bottom left) and Callisto (bottom right) datasets as determined from the association performed with the simulator. Both time-series and periodograms are shown. The lines indicate the position of the expected frequencies.
In this case the peaks of each satellite emerge without ambiguity at the expected position. Only for Io there is a less-significant, but still important, peak at about 1000 rad/y that had already been observed in the inclusive analysis of Fig. 5. In general the smaller peaks do not occur in correspondence of the expectations for other satellites supporting the fact that the rate of wrong-association is likely very low.
3.6 Sinusoidal fits and orbital parameters extraction
The elongation datasets, ($x_{i}$, $t_{i}$), separated by satellite have been fitted with a function of the form $x(t)=A\sin(\omega(t-t_{0})+\phi_{0})$ using the ROOT [6] fitter libraries. We have set the uncertainty on the point to match the r.m.s. of the residuals in practice setting the to the errors determined by the fit itself. The data with the superimposed best fita are shown in Fig. 8.
Figure 8: Fitted data points shown separately for each dataset associated to individual satellites. From top to bottom Io, Europa, Ganymede and Callisto. The units on the $x$-axis are fraction of years starting from 1/1/1610 and in the $y$-axis the elongation of the satellites in GJD units.
The same results are shown all together in Fig. 9.
Figure 9: Fitted data points shown separately for each dataset associated to individual satellites: Io (red), Europa (gray), Ganymede (orange) and Callisto (magenta). The units on the $x$-axis are fraction of years starting from 1/1/1610 and in the $y$-axis the elongation of the satellites in GJD units.
A video showing the result of the sinusoidal fits and the satellites’ associations is available in [7]. There the filled points show the prediction of the fit. At the time of observation the animation is briefly pause and the measurements from Galilei are superimposed together with an image of the original sketch.
The residuals of the fit ($x_{i}-x(t_{i})$) are shown for each satellite in Fig. 10. The top row shows them as a function of the point time while the bottom rows shows their distribution fitted with a Gaussian model.
Figure 10: Residuals for sinusoidal fits for Io (red), Europa (gray), Ganymede (orange) and Callisto (magenta). In the top row the units on the $x$-axis are fraction of years starting from 1/1/1610 and in the $y$-axis the elongation of the satellites in GJD units.
The model is reproducing the data very well without particular trends in the residuals or anomalies. In terms of the Jupiter disk in Galilei’s sketches (GJD) The r.m.s. of the residuals is 0.5 GJD for Io, 0.4 GJD for Europa, 0.56 GJD for Ganymede and 0.63 GJD for Callisto. Considering an average angular disk size of 42” during the measurements, and a factor 2.5 between the real and reported angular size, the previous r.m.s. correspond to values ranging from 40” to about 70”. This can be considered as an estimate of the pointing accuracy of the observations.
In the following we try to imagine how these data can be used to test the validity of Kepler third law for the Jupiter system and the 1:2:4 Laplace resonance between the periods of Io, Europa and Ganymede.
To do this test we have converted the radiuses of satellites’ orbits with a scaling factor chosen such that the scaled value for Ganymede corresponds to the modern measured value of 1.0704 million of km [^7] The scaling factor is 1.0704 Gm/$(4.482\pm 0.084)$ GJD = 2.39$\times$108 m/GJD..
Figure 11 and Tab. 7 show a comparison between the values of the satellites’ orbits radiuses $A$ and their periods $T=2\pi/\omega$ as found by Galilei (bullets) and by modern determinations (histogram).
quantity Galilei rel. err. (%) stat.
$T$ (Io) (1.7730$\pm$ 0.0018) days 0.10
$T$ (Europa) (3.5539 $\pm$ 0.0037) days 0.10
$T$ (Ganymede) (7.151$\pm$0.012) days 0.16
$T$ (Callisto) (16.678 $\pm$ 0.051) days 0.31
$A$ (Io) (0.443 $\pm$ 0.019) Gm 4.3
$A$ (Europa) (0.687 $\pm$ 0.012) Gm 1.7
$A$ (Ganymede) (1.070 $\pm$ 0.020) Gm 1.9
$A$ (Callisto) (1.6482 $\pm$ 0.020) Gm 1.9
Table 7: Comparison of $T$ and $A$ obtained with this analysis and the modern determinations. Distances are in units of million km and times in days.
Figure 11: Top left: satellites periods as derived from the sinusoidal fit to Galilei’s sketches (bullets) superimposed to the modern values (histogram). Top right: same as top left but for the major semi-axes. Bottom left: ratios between Galilei and modern for satellites’ periods. Bottom right: same as bottom left but for major semi-axes.
On the $x$-axis we show the satellite identifier (from left to right Io, Europa, Ganymede, Callisto). Periods are in seconds, are shown in the top left plot while radiuses (in meters) in the top right plot. The bottom plots show the ratio between the measured value by Galilei and the modern value. The periods are compatible with modern values within their errors that range from 0.1% to 0.3%. As far as the radiuses are concerned the ratios are compatible with unity with 1-2 $\sigma$ for Io and Europa. The one of Ganymede is one by construction while for Callisto $A$ is underestimated by about 12%. The statistical errors on these parameters range from 1.9% for the outer satellites to 4.3% for Io. There is hence a trend for a significant underestimation of the major semi-axes for Callisto (See Fig. 12, bottom right plot).
Figure 12 (left) shows a scatter plot of $A$ vs $T$. Superimposed a curve of the type $T^{\frac{2}{3}}$. The curve is the function: $A(T)=k_{J}T^{\frac{2}{3}}=\left(\frac{GM_{J}}{4\pi^{2}}\right)^{\frac{1}{3}}T^{\frac{2}{3}}$ with $G=6.67430\times 10^{-11}$ Nm2kg-2 and $M_{J}=1.898\times 10^{27}$ kg hence $A(T)=5.02\times 10^{5}\times T^{\frac{2}{3}}$.
Figure 12: Kepler third law for the Jupiter system.
In determining the law in physical units we have used our knowledge on the orbit of Ganymede to pass from angles to distances. It is interesting to notice that, if you use the knowledge on the distance of the Earth and Jupiter from the Sun, this measurement allows determining the ratio between the mass of Jupiter and that of the Sun just by comparing $k_{J}$ and $k_{Sun}$. Still you have to know that the Kepler constants depend on the mass to one third which is something that will not be clear until Newton’s law.
In Fig. 13 we compute the ratios $T_{Eu}/T_{Io}-2$ and $T_{Eu}/T_{Io}-4$ to verify the level of accuracy in determining the orbital resonance effects. Black filled bullets indicate the measurements of Galilei and the blue hollow ones the modern values. Good agreement is found. Deviations from an exact 1:2:4 ratios, also present in modern data, are due to the slow precession of Io’s orbit’s closest point to Jupiter (perijove).
Figure 13: Test of the 1:2:4 Laplace resonance for Io, Europa and Callisto.
3.7 The performances of the Sidereus Nuncius “cannocchiale”
Unfortunately the instrument used in early 1610 has been lost but we have other instruments that were used by Galilei. Some optical tests performed on two of these instruments is available at [8] and considerations on its aberrations in [9]. In this section we want to imagine what could have been the performances of the telescope based on the accuracy obtained on the Jupiter system. In the following we use sharp images and apply a very simple Gaussian smearing plus a deformation that brings objects close to the edge of the field of view to appear closer to the center than they really are. This is inspired by the observation that orbits of Ganymede and Callisto appear to be underestimated. This should be verified with an optical simulation or with a system that tries to reproduce the original devices. Furthermore the effect could be also simply related to the increased difficulty of correctly estimating the angular separation when the two points have a larger separation separated. Also it is likely that the distances were estimated in steps, and that Jupiter might have not always been in the center of the field. Despite of this we have nevertheless included this effect in the simulation. The idea is to try and get an impression of what would have been the image available to Galilei.
In Fig. 14 we show how the two effects act on the Jupiter system.
Figure 14: Top view is the ideal image of the Jupiter system. Center picture introduces spherical aberration while the bottom one also adds in the effect of resolution.
Applying the same smearing and aberration effects to terrestrial targets allows to get an insight of the impression that the observation with the “cannocchiale” might have given to people of those times. We know that in August 1609, i.e. about five months before the observations of Jupiter, Galilei had demonstrated the potentials of the cannocchiale to the Doge of Venice, Leonardo Donà (Fig. 15, left), probably with a telescope that was still not as good as the later one.
Here we assume that the observations were done from the Campanile of San Marco and, for simplicity, we use the performances of the telescope of January 1610. As a target we have chosen a quite peculiar on: the Scala Contarini del Bovolo that is a very distinctive building located at 360 m East of the observation point and with a diameter of about 6 m (Fig. 15, center and right). It would have surely attracted the attention of the viewers from the Campanile.
Figure 15: Left: Galileo and the Doge (Bertini’s fresco). Right: Scala Contarini del bovolo from the Campanile of San Marco. Campanile of San Marco.
The close up on the right of Fig. 15 is given to show the, very narrow, field of view of the Galileian telescope (circle) that we estimate in about 20” or a third of degree.
The view of the scala is shown in a modern detailed picture in Fig. 16 (left) while the other pictures show the effect of resolution (center) and resolution and aberration (right).
Figure 16: Simulation of the possible view of the Cannocchiale when watching scala Contarini del Bovolo from the Campanile of San Marco. The picture on the left is at high resolution. The middle one reproduces the limited resolution while the right one, both resolution and spherical aberration. The same smearing and aberration parameters used in Fig. 14 have been applied.
Despite the distortions and the narrow field of view it is clear how amazing this could have been and indeed we know how enthusiastically this technological advance was received, granting Galilei an immediate increase in salary.
4 Other observations in the Sidereus Nuncius
In Fig. 17, left, we show the drawing of Galilei of the iconic Pleiades star open cluster in the constellation of the Bull. On the right a modern image. The two patterns has been superimposed in the bottom part where we have attempted to maximize the overlap by tuning the relative scaling and rotation. The cluster extends for about 1.5 degrees while the field of view of the cannocchiale was likely about one fifth of that. This can explain why the two patterns show several systematic distortions as it could not be visible all together at once. Nevertheless it is easy to establish an unambiguous correspondence with the stars in the modern picture in almost all cases. The names and magnitudes of some of these stars are shown in the figure. Galileo could see stars up to magnitude almost 9 (i.e. 8.94 for HD2336). It is interesting to ask if the positions of some of the stars might really have changed in 400 years such that some differences might be real. Considering that the proper motions can arrive to 60 mas/year this would translate into an angular displacement of 24 as or about half a prime corresponding to about the size of he Jupiter disk. The contribution of this effect is hence very marginal.
Figure 17: Pleiades observation in the Sidereus Nuncius and comparison with modern images.
Galilei also reports observations of the Orion constellation. He writes that originally he wanted to accurately map the entire constellation but the task was overwhelming due to the large amount of visible stars. He then reports the region of the Orion belt where, as for the Pleiades, he marks the “old” (visible by naked eye) stars with large markers. A comparison of his map and the one from Stellarium is visible in Fig. 18.
Figure 18: The Orion belt region.
These maps resemble a bit old topographic maps of the World where basic features are correct but a lot of deformations are visible. We can state that with respect to the Pleiades the correspondence is less accurate. It should be noted that the field of view of the cannocchiale (about 1/3 of a degree) is tiny with respect to the extension of the map.
An observation of two “Nebulae” is also reported: one that Galilei calls the “Orion head nebula” and the Praesepe (nativity scene). This is also known as the Beehive cluster (Messier 44) in the Cancer constellation.
It is not clear what the head of Orion nebula might be. The natural expectation is that it refers to M42 whose nebular appearance is also at reach with naked eye, but it does not sit where the head of the Orion constellation is located. We have compared the sketch of Galilei with the group of stars located where the head of Orion sits but it seems unlikely that the sketch represents this region of the sky. More likely it shows the region just north of M42 as in Fig. 19. In this case the association is tentative and this should be considered as a reasonable guess. The stars within M42 also does not seem to match in a convincing way so the previous interpretation seems the most likely one.
Figure 19: The “head of Orion” nebula.
The comparison for the Beehive cluster or M44 is shown in Fig. 20. Also in this case the match is not evident so we present two possible interpretations. Galilei states that the pair of large stars are the two “little donkeys” (Asellos, currently known as $\gamma$ and $\delta$ Cancri) that are also visible by naked eye. In this case the comparison is the one of Fig. 20 a). Still it seems strange that stars are not so clustered in a small region as in reality. It could be related by the difficulty of representing accurately with the symbols the differences in brightness. It could even be that what Galilei represented is something more similar to the case b) i.e. that the two “little donkeys” might have in reality been other stars closer to the cluster. Also in this case it is rather hard to establish a one-to-one correspondence unlikely to the case of the Pleiades.
a)Refer to caption
b)Refer to caption
Figure 20: Beehive cluster comparison. Possible interpretations.
We also tried to compare the beautiful sketch of the first quarter Moon with some modern telescopic images. In this case the accuracy seems a bit worse than in the case of Jupiter. Galilei manages to convey very well the effect of shades at the terminator for craters and mountain ridges but the overall fidelity of the representation in terms of dimensions and arrangement is not at the same level. The crater in the bottom half along the terminator has a significantly overestimated size. It might correspond to the the big crater Ptolemaeus as Tycho is smaller and more south. The big mountain arc could correspond to the Alps-Caucasus-Apennines mountains which are located between the Sea of Serenity and the Sea of Showers. In the sketch of Galilei one can recognize the Sea pattern given by Sea of Serenity, Tranquillity, Fecundity and Crises and Nectar. The mountain arc seems to surround the Sea of Serenity instead of being in between the Sea of Serenity and the Sea of Showers as in reality.
Figure 21: Comparison of the Moon sketches with modern pictures.
5 Conclusions
In this work we have analyse the observations of the satellites of Jupiter reported by Galileo Galilei in the Sidereus Nuncius [2] during the period from January 7 to March 1, 1610.
We compare them to the predictions obtained using a modern sky simulator [3], verifying them one by one. A sinusoidal fit of the data obtained from the 64 available sketches, allows measuring the relative major semi-axes of the satellites’ orbits and their periods with a statistical precision of 2-4% and 0.1-0.3% respectively.
The periods are basically unbiased while the orbits tend to be underestimated for Callisto by about 12%. The posterior fit error indicates that the positions of the satellites are determined with a resolution of 0.4-0.6 Jupiter diameters in the notation of Galilei corresponding to about 40-70 arc sec i.e. similar to the true angular diameter of Jupiter, in those days.
We show that with this data one can infer in a convincing way the third law of Kepler for the Jupiter system. The 1:2 and 1:4 Laplace orbital resonance between the periods of Io and Europa/Ganymede can be determined with % precision.
In order to obtain these results it is important to separate the four datasets. This operation, which is nowadays simple using a sky simulator, and is fully reported in this work, was an extremely difficult task for Galilei as the analysis will evidence. Nevertheless we show how the four periods might have been extracted using the modern Lomb-Scargle technique without having to separate the four data-sets already just using these early observations.
We also perform a critical evaluation of the accuracy of the observation of the Pleiades cluster, the beehive cluster, the Orion belt, the Orion’s head nebula and the Moon.
To conclude we have also used these data to infer the possible resolution and aberration of the “cannocchiale” of Galilei and use these parameters to imagine what could have been the view offered by Galilei to the head of the Serenissima republic (the “Doge”) when he was promoting his instrument from the Campanile of San Marco in Venice.
This analysis allows quantifying and appreciating the level of accuracy that Galilei managed to attain with the instrumentation he had at his disposal. Overall the accuracy of the measurements is quite impressing and adds to the value of a giant of early experimental physics.
Appendix A Comparison of sketches and simulator
Figure 22: 1.
Figure 23: 2.
Figure 24: 3.
Figure 25: 4.
Figure 26: 5.
Figure 27: 6.
Figure 28: 7.
Figure 29: 8.
Figure 30: 9.
Figure 31: 10.
Figure 32: 11.
Figure 33: 12.
Figure 34: 13.
Figure 35: 14.
Figure 36: 15.
Figure 37: 16.
Figure 38: 17.
Figure 39: 18.
Figure 40: 19.
Figure 41: 20.
Figure 42: 21
Figure 43: 22
Figure 44: 23
Figure 45: 24
Figure 46: 25
Figure 47: 26
Figure 48: 27
Figure 49: 28
Figure 50: 29
Figure 51: 30
Figure 52: 31
Figure 53: 32
Figure 54: 33
Figure 55: 34
Figure 56: 35
Figure 57: 36
Figure 58: 37
Figure 59: 38
Figure 60: 39
Figure 61: 40
Figure 62: 41
Figure 63: 42
Figure 64: 43
Figure 65: 44
Figure 66: 45
Figure 67: 46
Figure 68: 47
Figure 69: 48
Figure 70: 49
Figure 71: 50
Figure 72: 51
Figure 73: 52
Figure 74: 53
Figure 75: 54
Figure 76: 55
Figure 77: 56
Figure 78: 57
Figure 79: 58
Figure 80: 59
Figure 81: 60
Figure 82: 61
Figure 83: 62
Figure 84: 63
Figure 85: 64
Appendix B Sunset correction
Figure 86: Sunset times in Padova as a function of the day since the beginning of the year. The red line shows a linear approximation that is used in the analysis to convert Galileo’s time data to the modern system as in Tab.  LABEL:tab:times.